OpenStudy (anonymous):

what is -6(2x+1)=4x+10

hero (hero):

Hint: 1. Use Distributive Rule: a(b+c) = ab + ac In this case: -6(2x + 1) = -12x - 6

OpenStudy (anonymous):

okay.

hero (hero):

Let me know what you get for x

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

is it - 1/2

OpenStudy (anonymous):

you solved -12x -6 = 0 great... but this problem is -12x -6 = 4x + 10

OpenStudy (anonymous):

can you help me solve the rest

OpenStudy (anonymous):

you do it that same way... just collect the terms get all the Xs on one side and everything else on the other

OpenStudy (anonymous):

add 12x to both sides and see what happens

hero (hero):

@philips13 , I know what the problem is. I was helping her with the first step, which is distribution.

OpenStudy (anonymous):

i solved it and i got -1 for x

OpenStudy (anonymous):

good job